# SqueezeSAM: User-Friendly Mobile Interactive Segmentation

### v1

@@ Abstract
Segment Anything Model (SAM) is a foundation model for interactive segmentation, and it has catalyzed major advances in generative AI, computational photography, and medical imaging. This model takes in an arbitrary user input and provides segmentation masks of the corresponding objects. It is our goal to develop a version of SAM that is appropriate for use in a photography app. The original SAM model has a few challenges in this setting. First, original SAM a 600 million parameter based on ViT-H, and its high computational cost and large model size that are not suitable for today’s mobile hardware. We address this by proposing the SqueezeSAM model architecture, which is 50x faster and 100x smaller than SAM. Next, when a user takes a photo on their phone, it might not occur to them to click on the image and get a mask. Our solution is to use salient object detection to generate the first few clicks; this produces an initial segmentation mask that the user can interactively edit. Finally, when a user clicks on an object, they typically expect all related pieces of the object to be segmented. For instance, if a user clicks on a person’s t-shirt in a photo, they expect the whole person to be segmented, but SAM typically segments just the t-shirt. We address this with a new data augmentation scheme, and the end result is that if the user clicks on a person holding a basketball, the person and the basketball are all segmented together.
1 Intro
Since the release of Segment Anything Model (SAM) [15], it has enabled leaps forward in generative AI, medical imaging, and computational photography [33, 27, 45]. Of particular interest to us is interactive segmentation for allowing users to select and extract objects within images on a smartphone. This functionality has already deployed in the iPhone with iOS 16, but the model is proprietary to Apple and cannot be retrained by researchers. Our goal is to create an open-source, fast, user-friendly, mobile model for interactive image segmentation. In our experience, one limitation of today’s mobile interactive segmentation interfaces is that many users will never click on their images, so they won’t discover the segmentation capability. This is unfortunate, because SAM expects the user to “make the first move" and click on an object to begin segmentation. To address this, we explore using salient object detection (SOD) to select a few interesting points in the image, which enables SAM to produce an initial segmentation that the user can refine by clicking on the image. One of the goals of SOD is to identify objects and regions where human eyes gravitate and fixate. So, we think it makes sense that SOD can guess some of the image regions that humans are likely to click on.
@@ 1 Intro
Another limitation of SAM is that it takes half a second on an A100 GPU and over ten seconds on iPhone CPU. Running locally on the device is important for interactive latency and for preserving privacy. The latency of SAM is dominated by its encoder, a ViT-H vision transformer [8]. Others have developed more efficient variants of SAM, but with dramatic reduction in the quality-of-results [3, 47, 48, 1]. In our work, we explore model architectures and training schemes that deliver low latency while producing quality-of-results that is more competitive with the original SAM model. Furthermore we find that our models can provide higher quality for segmenting salient objects.
@@
In summary, our novelties are a new model architecture that is 20x faster and 60x smaller than original SAM and an ability to produce segmentation masks for salient objects without user inputs. We will also publicly release the weights of our tiny model trained on 1 billion masks and 11 million images.
@@
Segment Anything Model (SAM) produced a dramatic improvement to interactive segmentation [15]. Similar to SimpleClick, SAM uses a ViT encoder. However, while SimpleClick uses a feature pyramid network decoder, the SAM decoder is Transformer-based. SAM’s most dramatic departure from the prior work is that it trained on a novel dataset of 11 million images and 1 billion masks. The new dataset, SA-1B, was a key ingredient in SAM’s ability to outperform FocalClick, SimpleClick, and RITM on interactive segmentation and instance segmentation benchmarks.
@@
[…]
Proposed approach
@@ 1.1 SqueezeSAM Encoder-Decoder Architecture
Our initial exploration began with using a VIT tiny backbone (as an alternative to VIT-H) trained on the SA-1B dataset. Due to the difficulty in quantization and deployment of VIT tiny models, subsequently, we also explored a UNet based architecture and discovered that we can train efficient and high quality models easily using a UNet backbone without much architecture tuning. We show our end-to-end model architecture in Figure 3. The inputs to our model are a image $\mathcal{I}\in\mathbb{R}^{3\times 1024\times 1024}$ image and a set of clicks produced by a user. [^2] For now, assume the clicks are from a user clicking a mouse or tapping a smartphone screen. In Section Application, we will explain how to use salient object detection to synthesize clicks. The model’s computational footprint is dominated by the SqueezeSAM encoder and decoder. In Figure 4, we show the precise layers and dimensions of our encoder and decoder, and we show the details of the DoubleConv blocks in Figures 56. Here are a few key design choices we made and why we made them:
@@ 1.3 Training
To achieve this relatively short training time on a large dataset, we changed how we feed clicks into the model during training. In the original SAM training, for each batch, the model runs on one click, then an click point is sampled and the model runs again. So, original SAM runs multiple steps of training for each batch of data. This inflates the training time, so for SqueezeSAM we simply input a collection of points for each batch and run one training step for each batch. Otherwise, we adhere to the original training protocol from the SAM paper.
Evaluation
@@
SqueezeSAM (Ours) 74.9 40.60 56.43 43.01 25.67 43.8 74.1 37.54 56.7 46.55 27.4 10 300 12.5
The disk size of our model is reduced to 12.5MB with int8 quantization. We find that with our architecture, we find minimal drop in model quality ( less than 1 percent).
In Table 1, we compare original SAM, SqueezeSAM, and MobileSAM. We observe that, relative to the other efficient models, SqueezeSAM produces higher-quality results. Our model is a UNET architecture that is trained from scratch on the SA-1B dataset. Unlike most other techniques that use tricks like pre-training and distillation, our models can get competitive performance without any of these techniques. We also find that our model has smaller latency compared to the mobileSAM version. In the table, note that SAM, MobileSAM, and SqueezeSAM were computed by the authors of this paper. For FastSAM, we use the numbers reported by the FastSAM authors.
[…]
@@
Application
Today’s mobile segmentation interfaces have a limitation: many users don’t click on images, missing out on the segmentation feature. To generate a few initial points to feed to SqueezeSAM, we considered “classic" (non-interactive) instance segmentation and also salient object detection (SOD). Qualitatively, we observed that SOD produced better input points that led to a superior interactive segmentation experience, so we used SOD. Another advantage of SOD is it can identify objects beyond the vocabulary of the training set, which usually includes multiple objects that are correlated in the context of the image, e.g. not just a person or a common object, but also a person holding a parrot.
The output of a SOD is a heatmap where the highest values of the heatmap refer to the most salient regions. To sample points from the saliency heatmap, after applying a global dynamic thresholding method like Otsu’s method [32], we detected the blobs. Most salient blob is identified with respect to the highest value within each blob, and we used the frequency of max heatmap value as a tie breaker when two blobs have the same value. After selecting the most salient blob, we divide it into 4 sections with respect the blob’s center of mass. Then we use the center of mass from each section as our selected points besides the center of mass of the whole blob, creating 5 clicks. When the center of mass lies out of the heatmap, we use the closest point within the mask. This way of sampling guarantees that the point distribution is in accordance to the blob shape. Fig shows a comparison to the grid sampling.
@@
Figure 8: From left to right: (1) A user clicks on a baseball player’s pants, (2) ground-truth mask, (3) original SAM output, (4) Salient SqueezeSAM output,. Notice that Salient SqueezeSAM segments the entire person, while other models only segment the pants.
Figure 9: From left to right: (1) A user clicks on a carrot, (2) ground-truth mask, (3) original SAM output, (4) Salient SqueezeSAM output. Notice that Salient SqueezeSAM segments the child holding the carrot, while other models just segment the carrot.
@@ Conclusions
Acknowledgements
We would like to thank Eric Mintun and Nikhila Ravi for providing valuable feedback during the model iteration. We would like to thank Bochuan Du, Terry Know, Chirag Ramanlal, Yang Qin and Avni Kakkar from the instagram team for providing valuable feedback on model quality when applied to instagram images. Finally would like to thank Lemeng Wu in help with creating visualizations and demo images.

### v3

@@ Abstract
The Segment Anything Model (SAM) has been a cornerstone in the field of interactive segmentation, propelling significant progress in generative AI, computational photography, and medical imaging. Despite its ability to process arbitrary user input and generate corresponding segmentation masks, SAM’s 600 million parameter architecture, based on ViT-H, is not compatible with current mobile hardware due to its high computational demands and large model size. Our research aims to adapt SAM for use in mobile photography applications. To this end, we have developed a fully convolutional SqueezeSAM model architecture, which is 62.5 times faster and 31.6 times smaller than the original SAM, making it a viable solution for mobile applications. Furthermore, our tiny model achieves an mIOU within 1% of the original VIT-H architecture.
Automated segmentation holds significant value in the creation flow for photography applications, as evidenced by its adoption by leading industry players like apple and capcut. To facilitate this automation, we employ salient object detection and simulate potential user clicks for foreground object selection, generating an initial segmentation mask that users can subsequently edit interactively. A common user expectation is that a click on a specific part of an object will result in the segmentation of the entire object. For example, a click on a person’s t-shirt in a photo should ideally segment the entire person, not just the t-shirt. However, SAM typically only segments the clicked area. We address this limitation through a novel data augmentation scheme. Consequently, if a user clicks on a person holding a basketball, both the person and the basketball are segmented together, aligning with user expectations and enhancing the overall user experience.
1 Introduction
Since the introduction of the Segment Anything Model (SAM) [15], it has catalyzed significant advancements in generative AI, medical imaging, and computational photography [33, 27, 45]. Our research focuses on interactive segmentation, a feature that enables users to select and extract objects within images on a smartphone. This functionality has already been implemented in the iPhone with iOS 16, but the proprietary model used by Apple is not available for retraining by researchers. Our objective is to develop an open-source, fast, user-friendly, mobile model for interactive image segmentation.
The adoption of automated segmentation by major industry players underscores its substantial importance for photography applications like apple and capcut. SAM requires the user to initiate the segmentation process by clicking on an object. To overcome this, we propose using salient object detection (SOD) to select a few interesting points in the image, enabling SAM to generate an initial segmentation that can be refined through user input. As SOD aims to identify objects and regions that attract human attention, it is logical to use it to predict the image regions that humans are likely to click on.
@@ 1 Introduction
Another challenge with SAM is its latency -it takes half a second on an A100 GPU and over ten seconds on an iPhone CPU. Local device operation is crucial for interactive latency and privacy preservation. The latency of SAM is primarily due to its encoder, a ViT-H vision transformer [8]. While more efficient SAM variants exist, they significantly compromise the quality-of-results [3, 47, 48, 1]. In our work, we explore model architectures and training schemes that offer low latency while maintaining a quality-of-results competitive with the original SAM model. We also show that our models deliver higher quality segmentation of salient objects.
@@
In summary, our novel contributions include a new model architecture that is 62.5 times faster and 31.6 times smaller than the original SAM, and the ability to generate segmentation masks for salient objects without user inputs. We will also publicly release the weights of our compact model, trained on 1 billion masks and 11 million images.
@@ 2 Related work
Segment Anything Model (SAM) has significantly advanced the field of interactive segmentation [15]. Like SimpleClick, SAM employs a ViT encoder. However, in contrast to SimpleClick’s feature pyramid network decoder, SAM utilizes a Transformer-based decoder. The most notable departure of SAM from previous work is its training on a novel dataset, SA-1B, comprising 11 million images and 1 billion masks. This new dataset was instrumental in enabling SAM to surpass FocalClick, SimpleClick, and RITM in both interactive and instance segmentation benchmarks.
@@
Salient Object Detection (SOD) involves predicting a heat map that distinguishes between salient and non-salient image regions.  [^1] To dispel a common point of confusion: While object detection typically means bounding boxes, SOD uses masks. Since 2010, several SOD benchmark datasets have emerged. In most benchmark datasets, humans label the ground-truth masks using a graphical interface [28, 37, 43]. Some datasets also incorporate human eye tracking data into the labeling process, based on the premise that areas of fixation are likely to be salient [18, 44]. Top performing models on SOD benchmarks include R3Net [7], U2Net [30], and SSOM [6]. Notably, SSOM employs SAM to enhance salient object detection. In a related vein, our research utilizes SOD to augment interactive segmentation.
3 Proposed approach
@@ 3.1 SqueezeSAM Encoder-Decoder Architecture
Our initial exploration began with using a VIT tiny backbone (as an alternative to VIT-H) trained on the SA-1B dataset. Due to the difficulty in quantization and deployment of VIT tiny models, subsequently, we also explored a UNet based architecture and discovered that we can train efficient and high quality models easily using a UNet backbone without much architecture tuning. We show our end-to-end model architecture in Figure 3. The inputs to our model are an image $\mathcal{I}\in\mathbb{R}^{3\times 1024\times 1024}$ and a set of clicks. [^2] For now, let’s assume the clicks are from a user clicking a mouse or tapping a smartphone screen. In Section 5, we will explain how to use salient object detection to synthesize clicks. The model’s computational footprint is dominated by the SqueezeSAM encoder and decoder. In Figure 4, we show the precise layers and dimensions of our encoder and decoder, and we show the details of the DoubleConv blocks in Figures 5 and 6. Here are a few key design choices we made and why we made them:
@@ 3.3 Training
To achieve this relatively short training time on a large dataset, we changed how we feed clicks into the model during training. In the original SAM training, for each batch, the model runs on one click, then another click point is sampled and the model runs again. So, original SAM runs multiple steps of training for each batch of data. This inflates the training time, so for SqueezeSAM we simply input a collection of points for each batch and run one training step for each batch. Otherwise, we adhere to the original training protocol from the SAM paper.
4 Evaluation
@@
EdgeSAM [50] 76.7 43.0 55.1 48.9 30.3 - 76.0 - - - - - - 9.6
EfficientSAM-Tiny [51] 75.7 42.3 57.4 46.2 26.7 74.3 39.9 59.9 51 28.9 19 - 10
EfficientSAM-Small [51] 77.0 44.4 60.1 48.3 28.4 75.4 42.3 62.3 54 30.8 21 - 25
SqueezeSAM (fp32) 77.9 41.1 57.0 45.1 25.1 44.5 78.1 37.4 58.0 49.6 25.5 8 300 19
SqueezeSAM (int8) 77.5 40.8 56.7 45.0 24.6 43.4 77.8 36.9 57.6 49.2 24.6 - 300 19
The disk size of our model is reduced to 14MB with int8 quantization. We find that with our architecture, we find less than 1 percent drop in model quality.
In Table 1, we compare original SAM, SqueezeSAM, and MobileSAM. We observe that, relative to the other efficient models, SqueezeSAM produces higher-quality results. Our model has a UNET architecture that is trained from scratch on the SA-1B dataset. Unlike most other techniques that use tricks like pre-training and distillation, our models can get competitive performance without any of these techniques. We also find that our model has smaller latency compared to the MobileSAM version.
(a) Input
(b) Saliency
(c) SAM [15]
(d) FastSAM [48]
(e) MobileSAM [47]
(f) SqueezeSAM (Ours)
@@
5 Application
The automatic generation of segmentation plays a crucial role in the development process for photography applications, as evidenced by its adoption by leading industry players such as Apple, CapCut. In line with these objectives, we explored ’classic’ (non-interactive) instance segmentation and salient object detection (SOD) to facilitate auto-creation. Our qualitative observations indicated that SOD helped us create superior input points, leading to an enhanced interactive segmentation experience. An additional advantage of SOD is its ability to identify objects beyond the vocabulary of the training set, often including multiple contextually correlated objects, such as a person holding a parrot, rather than just a person or a common object. Consequently, we opted to utilize SOD in our approach.
The output of a SOD is a heatmap where the highest values of the heatmap refer to the most salient regions. To sample points from the saliency heatmap, after applying a global dynamic thresholding method like Otsu’s method [32], we detected the blobs. Most salient blob is identified with respect to the highest value within each blob, and we used the frequency of max heatmap value as a tie breaker when two blobs have the same value. After selecting the most salient blob, we divide it into 4 sections with respect the blob’s center of mass. Then we use the center of mass from each section as our selected points besides the center of mass of the whole blob, creating 5 clicks. When the center of mass lies out of the heatmap, we use the closest point within the mask. This way of sampling guarantees that the point distribution is in accordance to the blob shape. Figure 8 shows a comparison to the grid sampling.
@@ 5 Application
Figure 9: From left to right: (1) User-click on original image (red-dot), (2) ground-truth mask from LVIS, (3) original SAM output, (4) Salient-SqueezeSAM output. Notice that Salient-SqueezeSAM segments the all correlated objects, while other models only segment single object instances.